### v1

@@ 7 Limitation and Future Work
With the repeated application of this equality, analogous to the above argument, we obtain $V_{B}\stackrel{{\scriptstyle s}}{{=}}V_{B}^{R_{A}(V)g^{*}}$. ∎
Appendix B Proof of Theorem 5.3
Proof.
First, we prove that $\hat{T}_{A}$ is a contraction mapping; hence it has a unique fixed point. For an arbitrary $(v,u)\in\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$ and $s\in\mathcal{S}$, we obtain
$(\hat{T}_{A}v)(s)-(\hat{T}_{A}u)(s)=\sup_{f\in\mathcal{W}_{A}}\left\{r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[v(s^{\prime})\right]\right\}\\ -\sup_{f\in\mathcal{W}_{A}}\left\{r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[u(s^{\prime})\right]\right\}$ (20)
from the definition. Let the first term on the right-hand side of the above equation be $v^{*}(s)$. Then, for an arbitrary $\delta>0$, there exists $f\in\mathcal{W}_{A}$ such that
$r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[v(s^{\prime})\right]\geq v^{*}(s)-\delta.$ (21)
Let $\delta=\frac{1-\gamma_{A}}{2}\lVert v(s)-u(s)\rVert_{\infty}$ and $f_{v}$ satisfy the above condition. Then,
$(\hat{T}_{A}v)(s)-(\hat{T}_{A}u)(s)$ $\leq v^{*}(s)-\left\{r_{A}(s,f_{v}(s),R_{B}^{*}(s,f_{v}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{v}(s),R_{B}^{*}(s,f_{v}))}\left[u(s^{\prime})\right]\right\}$ (22a)
[…]
$=\delta+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{v}(s),R_{B}^{*}(s,f_{v}))}\left[v(s^{\prime})-u(s^{\prime})\right]$ (22e)
$\leq\delta+\gamma_{A}\lVert v(s)-u(s)\rVert_{\infty}$ (22f)
$=\frac{1+\gamma_{A}}{2}\lVert v(s)-u(s)\rVert_{\infty}.$ (22g)
Analogously, we have
$(\hat{T}_{A}u)(s)-(\hat{T}_{A}v)(s)\leq\frac{1+\gamma_{A}}{2}\lVert v(s)-u(s)\rVert_{\infty}.$ (23)
Therefore, since $s$ is arbitrary, we have
$\lVert(\hat{T}_{A}v)(s)-(\hat{T}_{A}u)(s)\rVert_{\infty}\leq\frac{1-\gamma_{A}}{2}\lVert v(s)-u(s)\rVert_{\infty}$ (24)
for any $(v,u)\in\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$. This implies that $\hat{T}_{A}$ is a contraction mapping. According to the Banach fixed-point theorem, $\hat{T}_{A}$ has a unique fixed point.
Subsequently, we prove that $\hat{V}_{A}^{*}$ is the fixed point of $\hat{T}_{A}$. From the definition, we obtain
$\hat{V}_{A}^{*}(s)$ $=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)$ (25a)
$=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[V_{A}^{\hat{\mathbf{f}}{\dagger}}(s^{\prime})\right]$ (25b)
$\leq\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (25c)
$=\sup_{\hat{f}_{s}\in\mathcal{W}_{A}}r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (25d)
$=(\hat{T}_{A}\hat{V}_{A}^{*})(s)$ (25e)
for all $s\in\mathcal{S}$. We show that the equality in Equation 25c holds for all $s\in\mathcal{S}$. Suppose that there exists $s_{0}\in\mathcal{S}$ such that the above inequality strictly holds. Let the discrepancy between the left-hand side and the right-hand side of the above inequality at state $s_{0}$ be $2\epsilon$. Then, there must exists $\hat{\mathbf{f}}\in\mathcal{W}_{\mathcal{A}}^{\mathcal{S}}$ such that
$\hat{V}_{A}^{*}(s)$ $\leq r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]-\epsilon\quad\text{for }s=s_{0}$ (26)
$\hat{V}_{A}^{*}(s)$ $\leq r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]+\epsilon(1-\gamma_{A})\quad\text{for }s\neq s_{0}$ (27)
holds, where $\hat{f}_{s}$ is the policy derived from $\hat{\mathbf{f}}$ at $s$. Then, starting $s_{0}$ with Equation 26 and repeatedly applying Equation 27, we obtain
$\hat{V}_{A}^{*}(s_{0})$
$\leq r_{A}(s_{0},\hat{f}_{s_{0}}(s_{0}),R_{B}(s_{0},\hat{f}_{s_{0}})+\gamma_{A}\mathbb{E}_{s_{1}}\left[\hat{V}_{A}^{*}(s_{1})\right]-\epsilon$ (28a)
$\end{multlined}r_{A}(s_{0},\hat{f}_{s_{0}}(s_{0}),R_{B}(s_{0},\hat{f}_{s_{0}}))+\gamma_{A}\mathbb{E}_{s_{1}}\Bigl{[}r_{A}(s_{1},\hat{f}_{s_{1}}(s_{1}),R_{B}(s_{1},\hat{f}_{s_{1}}))+\gamma_{A}\mathbb{E}_{s_{2}}\left[\hat{V}_{A}^{*}(s_{2})\right]\Bigl{]}-\epsilon+\epsilon\gamma_{A}(1-\gamma_{A})$ (28c)
[…]
$\leq\mathbb{E}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}\left[\sum_{t=0}^{\infty}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right]-\epsilon+\epsilon(1-\gamma_{A})\sum_{t=0}^{\infty}\gamma_{A}^{t+1}$ (28d)
$=V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)-\epsilon(1-\gamma_{A}).$ (28e)
That is, $V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)>\hat{V}_{A}^{*}(s_{0})$, contradicting $\hat{V}_{A}^{*}\stackrel{{\scriptstyle s}}{{=}}\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}$. Therefore, the equality must hold in Equation 25c for all $s\in\mathcal{S}$. This implies
$\hat{V}_{A}^{*}(s)=\max_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]=(\hat{T}_{A}\hat{V}_{A}^{*})(s).$ (29)
This completes the proof. ∎
Appendix C Proof of Lemma 5.4
Proof.
Let $\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}$ satisfy $\hat{f}_{s}\in R_{A}^{*}(s)$ for all $s\in\mathcal{S}$. When the follower follows the best response against $\hat{\mathbf{f}}$, from the definition of $R_{A}^{*}$, we have
$r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]\\ =\max_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (30)
for all $s\in\mathcal{S}$. Therefore, from Equation 29, it holds that
$\hat{V}_{A}^{*}(s)=r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (31)
for all $s\in\mathcal{S}$. Let an operator $T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}:\mathcal{F}_{\mathcal{S}}\to\mathcal{F}_{\mathcal{S}}$ be defined as
$(T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}v)(s)=r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[v(s^{\prime})\right],$ (32)
then $T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}$ is a contraction mapping, and as it holds that
$V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}(s)$ $=r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}(s^{\prime})\right]$ (33a)
$=(T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})})(s)$ (33b)
for all $s\in\mathcal{S}$ from the definition of $V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}$, $V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}$ is the unique fixed point of $T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}$.
From Equation 31, since $\hat{V}_{A}^{*}$ satisfies $\hat{V}_{A}^{*}(s)=(T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}\hat{V}_{A}^{*})(s)$ for all $s\in\mathcal{S}$, it holds that $V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)=\hat{V}_{A}^{*}(s)$ for all $s\in\mathcal{S}$. ∎
Appendix D Proof of Theorem 5.5
Proof.
From Lemma 5.4, it follows that if there exists $f^{\prime}\in\mathcal{W}_{A}$ such that $\hat{f}\in\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$, it holds that $V_{A}^{\hat{f}{\dagger}}(s)=\hat{V}_{A}^{*}(s)$ for all $s\in\mathcal{S}$. Then, because it holds that
$\hat{V}_{A}^{*}(s)=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)\geq\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$ (34)
from $\mathcal{W}_{A}^{\mathcal{S}}\supset\mathcal{W}_{A}$, it holds that
$\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)\geq V_{A}^{\hat{f}{\dagger}}(s)=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)\geq\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$ (35)
for all $s\in\mathcal{S}$. Therefore, it holds that $V_{A}^{\hat{f}{\dagger}}(s)=\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$ for all $s\in\mathcal{S}$. ∎
Appendix E Proof of Corollary 5.6
Proof.
From Theorem 5.5, if $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ is non-empty, then a SSE policy exists.
If $\hat{f}\in\mathcal{W}_{A}$ is a SSE policy, it is holds that $\hat{f}\in\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$ for all $s\in\mathcal{S}$ from the definition. If $\gamma_{A}=0$, because
$R_{A}^{*}(s)=\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)=\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))$
holds, $\hat{f}\in R_{A}^{*}(s)$ for all $s\in\mathcal{S}$. It implies that $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ is nonempty. ∎
Appendix F Proof of Proposition 6.2
Proof.
@@
[…]
Altogether, we have $V_{A}^{*}\in\mathcal{PV}$ if and only if $\mathcal{PV}$ is a singleton. Because PO policies exists if and only if $\mathcal{PV}\cap\mathcal{V}\neq\emptyset$, the SE policy exists if and only if $\mathcal{PV}=\{v^{*}\}$ is a singleton and $v^{*}\in\mathcal{V}$. ∎
Appendix G Proof of Lemma 6.3
Proof.
@@
If it holds that $Q_{A}^{f^{\prime}{\dagger}}(s,f)=V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}$, by the similar derivation of Equation 37, we obtain
@@
The policy improvement is possible if and only if $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succ V_{A}^{f^{\prime}{\dagger}}$. Because of the above equality, equivalently we can say that the policy improvement from $f^{\prime}$ is possible if and only if there exists a policy $f$ whose value function satisfies
@@
for all $s\in\mathcal{S}$, and there exists a state $s$ where the inequality strictly holds. ∎
Appendix H Proof of Lemma 6.4
Proof.
We prove the counterpart of Lemma 6.4. Taking the negation of Equation 7 for all $f\in\mathcal{W}_{A}$, we have
@@
From Lemma 6.3 (a), since $V_{A}^{f{\dagger}}(s)\geq V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}$ holds and $Q_{A}^{f^{\prime}{\dagger}}(s,f)\geq V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}$ holds, we have
@@
Therefore, $f^{\prime}$ is not a PO policy. This completes the proof. ∎
Appendix I Proof of Lemma 6.5
Proof.
@@
$\geq r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\Bigl{(}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[Q_{A}^{f^{\prime}{\dagger}}(s^{\prime},f)\right]-\delta(f,f^{\prime})\Bigl{)}$ (50b)
@@
$\geq\mathbb{E}^{fR_{B}^{*}(f)}\left[\sum_{t=0}^{n-1}\gamma_{A}^{t}r_{A}(s_{t},f(s_{t}),R_{B}^{*}(s_{t},f))+\gamma_{A}^{n}V_{A}^{f^{\prime}{\dagger}}(s_{n})\middle|s_{0}=s\right]-\left(\sum_{t=0}^{n-1}\gamma_{A}^{t}-\gamma_{A}^{0}\right)\delta(f,f^{\prime}).$ (50c)
for all $s\in\mathcal{S}$. In the limit for $n\to\infty$, the first term converges to $V_{A}^{f{\dagger}}(s)$, the second term converges to 0, and the third term converges to $-\frac{\gamma_{A}}{1-\gamma_{A}}\delta(f,f^{\prime})$. Therefore, we have
@@
If (ii) holds, because of Lemma 6.3 (b), it holds that
@@
Therefore,
Considering this, it holds that
$\forall f\in\mathcal{W}_{A}\{\mathrm{(i)~{}or~{}(ii)}\}\implies\forall f\in\mathcal{W}_{A}\left\{V_{A}^{f^{\prime}{\dagger}}(s)>V_{A}^{f{\dagger}}(s)~{}~{}\exists s\in\mathcal{S}\mathrm{~{}~{}or~{}~{}}V_{A}^{f^{\prime}{\dagger}}(s)=V_{A}^{f{\dagger}}(s)~{}~{}\forall s\in\mathcal{S}\right\},$ (55)
where the right-hand side is equivalent to the definition of PO policies (Definition 6.1). ∎
Appendix J Proof of Theorem 6.6
Proof.
We prove each statement one by one.
@@
$\mathcal{L}\left[V_{A}^{f_{t+1}{\dagger}}\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\geq\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f_{t+1})\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\geq(1-\epsilon)\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\Bigl{(}\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\Bigl{)}.$ (57)
@@
therefore, for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$, we have $\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]=\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]$. Because $\mathcal{L}$ is Pareto-compliant, the above equality implies $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\not\succ V_{A}^{f_{t}{\dagger}}$. However, because $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$, we have $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f_{t}{\dagger}}$, implying $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$. From Statement (b) of Lemma 6.3, we have $V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$. (Conversely, if $V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$, we have $v_{t+1}\stackrel{{\scriptstyle s}}{{=}}v_{t}$.) Thus, $\mathcal{W}_{\succcurlyeq}(f)=\mathcal{W}_{\succcurlyeq}(f_{t})$ holds for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$. It implies that $\mathcal{W}_{\succcurlyeq}(f_{t+1})=\mathcal{W}_{\succcurlyeq}(f_{t})$ since $f_{t+1}\in\mathcal{W}_{\succcurlyeq}(f_{t})$; thus, $f_{t+2}\in\mathcal{W}_{\succcurlyeq}(f_{t+1})=\mathcal{W}_{\succcurlyeq}(f_{t})$. Therefore,
@@
First, we show that $\min_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\right)$ for all $v\in\mathcal{V}$.
Suppose that there exists $v^{*}\in\mathcal{V}$ such that $\min_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))>\gamma_{A}\left(\max_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))\right)$ holds. Then, $v^{*}\succ v_{\infty}$ and there exists $f^{*}\in\mathcal{W}_{A}$ whose value function is $V_{A}^{f^{*}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}v^{*}$. Let $\tau=\frac{\min_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))}{\max_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))}$. We have $1\geq\tau>\gamma_{A}$.
For an arbitrarily small $s>0$ there exists $t$ such that $\lVert v_{t}-v_{\infty}\rVert_{\infty}\leq s$ because $v_{\infty}\stackrel{{\scriptstyle s}}{{=}}\lim_{t\to\infty}v_{t}$. Because of the continuity of $\mathcal{L}$, it also implies that for an arbitrarily small $s^{\prime}>0$ there exists $t^{\prime}$ such that $\mathcal{L}[v_{t^{\prime}}]>\mathcal{L}[v_{\infty}]-s^{\prime}$. Therefore, we can find $t$ satisfying

### v2

@@ 7 Limitation and Future Work
With the repeated application of this equality, analogous to the above argument, we obtain $V_{B}\stackrel{{\scriptstyle s}}{{=}}V_{B}^{R_{A}(V)g^{*}}$.
Appendix B Proof of Proposition˜6.2
Proof B.1.
@@
Subsequently, suppose that $\mathcal{PV}=\{v^{*}\}$ is a singleton. Then, $v^{*}\succ v$ for all $v\in\mathcal{V}\setminus\mathcal{PV}$ because, for all $v\in\mathcal{V}\setminus\mathcal{PV}$, there exists $\hat{v}\in\mathcal{PV}$ such that $\hat{v}\succ v$. Suppose that $v^{*}\not\stackrel{{\scriptstyle s}}{{=}}V_{A}^{*}$. Then, for some $s\in\mathcal{S}$, $v^{*}(s)<\sup_{f\in\mathcal{W}}V_{A}^{f{\dagger}}(s)$ holds. It implies that there exists $v\in\mathcal{V}\setminus\mathcal{PV}$ such that $v^{*}(s)<v(s)$ for some $s$, which contradicts $v^{*}\succ v$. Hence, $v^{*}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{*}$.
Altogether, we have $V_{A}^{*}\in\mathcal{PV}$ if and only if $\mathcal{PV}$ is a singleton. Because PO policies exists if and only if $\mathcal{PV}\cap\mathcal{V}\neq\emptyset$, the SE policy exists if and only if $\mathcal{PV}=\{v^{*}\}$ is a singleton and $v^{*}\in\mathcal{V}$.
Appendix C Proof of Theorem˜6.3
Proof C.1.
@@
If it holds that $Q_{A}^{f^{\prime}{\dagger}}(s,f)=V_{A}^{f^{\prime}{\dagger}}(s)~\forall s\in\mathcal{S}$, by the similar derivation of Equation˜19, we obtain
@@
The policy improvement is possible if and only if $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succ V_{A}^{f^{\prime}{\dagger}}$. Because of the above equality, equivalently, we can say that the policy improvement from $f^{\prime}$ is possible if and only if there exists a policy $f$ whose value function satisfies
@@
for all $s\in\mathcal{S}$, and there exists a state $s$ where the inequality strictly holds.
Appendix D Proof of Theorem˜6.4
Proof D.1.
We prove the counterpart of Theorem˜6.4. Taking the negation of Equation˜5 for all $f\in\mathcal{W}_{A}$, we have
@@
From Theorem˜6.3 (a), since $V_{A}^{f{\dagger}}(s)\geq V_{A}^{f^{\prime}{\dagger}}(s)~\forall s\in\mathcal{S}$ holds and $Q_{A}^{f^{\prime}{\dagger}}(s,f)\geq V_{A}^{f^{\prime}{\dagger}}(s)~\forall s\in\mathcal{S}$ holds, we have
@@
Therefore, $f^{\prime}$ is not a PO policy. This completes the proof.
Appendix E Proof of Theorem˜6.5
Proof E.1.
@@
$\geq\dots$
[…]
for all $s\in\mathcal{S}$. In the limit for $n\to\infty$, the first term converges to $V_{A}^{f{\dagger}}(s)$, the second term converges to 0, and the third term converges to $\frac{\gamma_{A}}{1-\gamma_{A}}\delta(f,f^{\prime})$. Therefore, we have
@@
If (ii) holds, because of Theorem˜6.3 (b), it holds that
@@
Therefore, it holds that
$\forall f\in\mathcal{W}_{A}\{\mathrm{(i)~or~(ii)}\}\\ \implies\forall f\in\mathcal{W}_{A}\left\{V_{A}^{f^{\prime}{\dagger}}(s)>V_{A}^{f{\dagger}}(s)~~\exists s\in\mathcal{S}\mathrm{~~or~~}V_{A}^{f^{\prime}{\dagger}}(s)=V_{A}^{f{\dagger}}(s)~~\forall s\in\mathcal{S}\right\},$ (37)
where the right-hand side is equivalent to the definition of PO policies (Definition˜6.1).
Appendix F Proof of Theorem˜6.6
Proof F.1.
We prove each statement one by one. The proof of (a) is based on the monotonicity of $\{v_{t}\}$ and the compactness of $\mathrm{cl}\mathcal{V}$. The proof of (b) is based on the statement (b) of Theorem˜6.3. The statements (c) and (d) are proved by contradiction using the statement (c) of Theorem˜6.3.
@@
$\mathcal{L}\left[V_{A}^{f_{t+1}{\dagger}}\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]$ $\geq\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f_{t+1})\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]$ (39)
$\geq(1-\epsilon)\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\Bigl(\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\Bigl).$ (40)
@@
therefore, for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$, we have $\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]=\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]$. Because $\mathcal{L}$ is Pareto-compliant, the above equality implies $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\not\succ V_{A}^{f_{t}{\dagger}}$. However, because $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$, we have $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f_{t}{\dagger}}$, implying $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$. From Statement (b) of Theorem˜6.3, we have $V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$. (Conversely, if $V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$, we have $v_{t+1}\stackrel{{\scriptstyle s}}{{=}}v_{t}$.) Thus, $\mathcal{W}_{\succcurlyeq}(f)=\mathcal{W}_{\succcurlyeq}(f_{t})$ holds for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$. It implies that $\mathcal{W}_{\succcurlyeq}(f_{t+1})=\mathcal{W}_{\succcurlyeq}(f_{t})$ since $f_{t+1}\in\mathcal{W}_{\succcurlyeq}(f_{t})$; thus, $f_{t+2}\in\mathcal{W}_{\succcurlyeq}(f_{t+1})=\mathcal{W}_{\succcurlyeq}(f_{t})$. Therefore,
@@
First, we show that
$\min_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\leq\gamma_{A}(\max_{s\in\mathcal{S}}(v(s)-v_{\infty}(s)))$
for all $v\in\mathcal{V}$.
Suppose that there exists $v^{*}\in\mathcal{V}$ such that
$\min_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))>\gamma_{A}(\max_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s)))$
holds. Then, $v^{*}\succ v_{\infty}$ and there exists $f^{*}\in\mathcal{W}_{A}$ whose value function is $V_{A}^{f^{*}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}v^{*}$. Let $\tau=\frac{\min_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))}{\max_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))}$. We have $1\geq\tau>\gamma_{A}$.
For an arbitrarily small $\xi>0$ there exists $t$ such that $\lVert v_{t}-v_{\infty}\rVert_{\infty}\leq\xi$ because $v_{\infty}\stackrel{{\scriptstyle s}}{{=}}\lim_{t\to\infty}v_{t}$. Because of the continuity of $\mathcal{L}$, it also implies that for an arbitrarily small $\xi^{\prime}>0$ there exists $t^{\prime}$ such that $\mathcal{L}[v_{t^{\prime}}]>\mathcal{L}[v_{\infty}]-\xi^{\prime}$. Therefore, we can find $t$ satisfying